If three angles of a triangle are in the ratio of 1:2:5:.find their angles

Grade:7

3 Answers

Priyanka Chatterjee
7 years ago
Let the angles be x, 2x, 5xSo...x+2x+5x=1808x=180 x= 22.5It means the angles are 22.5°, 45° & 112.5°
Anand kumar
7 years ago
HiLet be ratio x the x+2x+5x=1808x=180 then x=22.5 then the angle 22.5°,2×22.5=45°and 5×22.5°=112.5°
Digvijay
6 years ago
Let angles be x, 2x, 5x.=x+2x+5x= 180 =8x=180=x=22.5Putting the values in the ratiosx=22.5°2x=45°5x=112.5°Ans= 22.5°, 45°, 112.5°